OpenStudy (anonymous):

Whitney's delivery truck has a weight limit of 1.5 tons. She delivers refrigerators that weigh 225 lb, dishwashers that weigh 150 lb, and ovens that weigh 300 lb. (Hint: 1 ton = 2000 lb.) a. Write a system of equations in three variables to represent the situation. b. Complete the table for the possible numbers of appliances the truck can hold. Refrigerators: 8, __, 4, 10 Dishwashers: 6, 1, __, 5 Ovens: __, 5, 4, __

OpenStudy (valpey):

Refrigerators: 8, x, 4, 10 Dishwashers: 6, 1, y, 5 Ovens: z, 5, 4, w Fridges are 225, DWs 150, Ov 300 1.5 tons = 1.5*2000lbs = 3000 lbs. = max weight 3000 >= 8*225 + 6*150 + z*300 3000 >= x*225 + 1*150 + 5*300 3000 >= 4*225 + y*150 + 4*300 3000 >= 10*225 + 5*150 + w*300

OpenStudy (valpey):

Just remember x, y, z, and w can only be non-negative integers.